Poiseuille Flow: Consider a pressure-driven (gradi
Answer #1

here,

4.

magnitude of charges , q = 4.65 * 10^-10 C

electric field , E = 1660 N/C

let the distance between the charges be x

as both are equal in magnitude

electric feild , E = k * q / (0.5 * x)^2 + k * q / (0.5 * x)^2

2 * 4.65 * 10^-10 * 9 * 10^9 / ( 0.5 x)^2 = 1660

x = 0.14 m

distance between the charrges is 14 cm or 0.14 m

5.

for the charges of equal sign,

as the charges are equal

the magnitude of electric feild at the centre is zero